### v3

Abstract
ZKP systems have surged attention and held a fundamental role in contemporary cryptography. Zk-SNARK protocols dominate the ZKP usage, often implemented through arithmetic circuit programming paradigm. However, underconstrained or overconstrained circuits may lead to bugs. Underconstrained circuits refer to circuits that lack the necessary constraints, resulting in unexpected solutions in the circuit and causing the verifier to accept a bogus witness. Overconstrained circuits refer to circuits that are constrained excessively, resulting in the circuit lacking necessary solutions and causing the verifier to accept no witness, rendering the circuit meaningless. This paper introduces a novel approach for pinpointing two distinct types of bugs in ZKP circuits. The method involves encoding the arithmetic circuit constraints to polynomial equation systems and solving polynomial equation systems over a finite field by algebraic computation. The classification of verification results is refined, greatly enhancing the expressive power of the system. We proposed a tool, AC4, to represent the implementation of this method. Experiments demonstrate that AC4 represents a substantial 29% increase in the checked ratio compared to prior work. Within a solvable range, the checking time of AC4 has also exhibited noticeable improvement, demonstrating a magnitude increase compared to previous efforts.
Keywords:
zero knowledge proofscircuit constraintsunderconstrainedoverconstrainedalgebraic computation checker
1 Introduction
Zero-knowledge proof (ZKP) systems have long occupied a fundamental position in contemporary cryptography, ever since their inception [21]. They are significantly utilized in security-sensitive applications, such as smart contracts [44], blockchain confidential transactions [44, 22], and digital currency [49], given their capacity to bolster security and privacy. ZKP systems’ potency lies in their ability to prove anything that an interactive proof system can, but without revealing any knowledge [24, 7]. Therefore, these systems have assumed the role of a potent tool enabling trust and authentication while safeguarding sensitive information.
Developers often opt for the zero-knowledge succinct noninteractive argument of knowledge (zk-SNARK) [9] as the protocol for the efficiency of setting up and proving of building a ZKP system, while also adopting arithmetic circuit [39] as the programming paradigm due to its ease of programming. Circom, a domain-specific language, is based on the aforementioned protocol and programming paradigm. It is capable of converting computational statements into constraints, which can be used to generate a zk system with snark tools such as snarkjs. To illustrate, consider the process of declaring a boolean value in circuit programs. In high-level languages, a boolean value $b$ can be declared directly. However, in circuit languages such as Circom, a constraint must be explicitly defined to limit its possible values, such as the constraint $b\times(b-1)=0$ for $b$. This transformation cannot be automatically performed by existing compilers or libraries. Thus, developers are compelled to express the computations with polynomial equations.
The limitation of polynomial constraint systems gives rise to two problems. The first problem is the underconstrained problem [37], which occurs when multiple outputs can fulfill the same system of equations with identical input values. This situation arises when the system of polynomial equations does not yield a mathematical function on the finite field, rendering the circuit underconstrained. As a result, malicious users could generate fake witnesses, deceive the verifier, and potentially steal the cryptocurrency [35, 14]. The second problem is the overconstrained problem [37], which occurs when no outputs can satisfy the system of polynomial equations or the constraint system is inconsistent for a given input. This problem may give rise to a useless circuit, which in turn opens up the possibility of a denial of service attack. It is obvious that the underconstrained circuits could bring more severe results than the overconstrained circuits.
Identifying underconstrained and overconstrained problems can be achieved by encoding their properties into a logical formula and utilizing a Satisfiability Modulo Theories (SMT) solver to solve them directly. Picus [37] is the first work that makes a valuable effort to employ an SMT solver for investigating the underconstrained cases of ZKP circuits with static program analysis. However, the overconstrained issue still lacks resolution. Nevertheless, the efficiency of SMT solvers brings an obstacle to SMT-based tools when solving large circuits. SMT-based tools struggle to verify constrained circuits within a limited time, even for linear ones.
To this end, an investigation of the constraints on Circom circuit programs is carried out. By reconsidering the formalization paradigm, we model underconstrained and overconstrained problems as algebraic inquiries into whether a quadratic equations system over a finite field has multiple solutions. An algebraic approach is a more intuitive and efficient way to proceed with the realization that the highest degree of the quadratic equations system of equations not exceeding two. Moreover, to support our motivation, we analyzed over 100 popular Circom circuit programs and observed that nearly 70% of them are linear while efficient algorithms exist for solving linear systems of equations. Therefore, after categorizing constraint problems into diverse types according to their maximal degree, we employ a unified arithmetic solution framework to solve them separately.
To summarize, contributions can be elaborated as follows:
  • Modeling Circom circuits as polynomial systems over finite fields, we categorize them and propose specialized methods for each category.
  • We develop $\text{AC}^{4}$, a complete and sound tool for verifying circuit programs with the framework above.
  • We evaluate the effectiveness through the solved ratio and solving speed of circuit programs. Our approach solved 65% of previously unsolved examples, a 29% increase over prior work, and significantly improved solving speed.
2 Background
Zk-SNARK, the dominant ZKP protocol in mainstream tools, frequently uses arithmetic circuits for programming. The verification efforts focus on Circom circuits, which leverage both zk-SNARK and arithmetic circuits.
Zk-SNARKs. Non-interactive ZKPs, such as zk-SNARKs [8], allow verifiers to be assured of a statement’s truth without gaining any knowledge beyond its validity [21]. Blum et al. demonstrated that computational zero-knowledge can be achieved without interaction by sharing a common random string between the prover and the verifier [13, 48]. The non-interactive protocol conveys the proof through a message from the prover to the verifier, ensuring efficiency. zk-SNARKs reign supreme in zk systems due to lightning-fast verification of even large programs, using tiny proofs (mere hundreds of bytes) within milliseconds.
Arithmetic Circuits. Arithmetic circuits, fundamental models for polynomial computation in complexity theory, process variables/numbers using addition and multiplication. Conceptually, arithmetic circuits offer a formal method for understanding the complexity of computing polynomials. Note that all variables mentioned in arithmetic circuits represent signals in the Circom language.
Definition 1 (Arithmetic circuit [45])
An arithmetic circuit $C(\mathcal{X})$ in the variables $\mathcal{X}=\{X_{1},\ldots,X_{k}\}$ over the field $\mathbb{F}$ is a labelled directed acyclic graph. The inputs are assigned by variables from $X$ or by constants. The internal nodes act as computational operations that add or multiply to calculate the sum or product of the polynomials on the tails of incoming edges. The output $X_{out}\in\mathcal{X}$ of a circuit is the polynomial computed at the output node.
Circom. Circom is a compiler written in Rust [33] for compiling circuits written in the Circom language, a SNARK-based language that uses arithmetic circuits as its programming paradigm. The compiler outputs circuit constraints and all the necessary information for computing different ZK proofs. In Circom, programmers can define constraints using the operators ======= = = for equality, <<--< - - or >-->- - > for signal assignment, and <==absent<==< = = or ==>==>= = > for simultaneous assignment. Constraints in Circom must be in the form of $A\times B+C=0$, where $A$, $B$, and $C$ are linear combinations of signals. After compiling circuit to r1cs format and computing the witness with the input, users can utilize the snarkjs tool to generate and validate a proof for the specified input. The verifier will receive the proof if the witness is valid, otherwise, the verifier will reject it.
We address Circom circuit constraint verification by extending arithmetic circuit definitions with constraints.
Definition 2 (Constraint)
A constraint $e$ is a quadratic equation over a finite field $F_{p}$:
$\mathcal{A}\mathcal{X}\times\mathcal{B}\mathcal{X}+\mathcal{C}\mathcal{X}=0$
where $\mathcal{A},\mathcal{B},\mathcal{C}$ are coefficient vectors over $F_{p}$, $\mathcal{X}$ is a vector of variables over $F_{p}$, and the cross product $\times$ means the product of 2 scalars $\pmod{p}$.
Definition 3 (Arithmetic circuit with constraints)
An arithmetic circuit with constraint with constraints can be defined as a tuple $\langle C(\mathcal{X}),E\rangle$, where $C(\mathcal{X})$ is an arithmetic circuit. $\mathcal{X}$ is split into two disjoint sets $K$ and $U$. $K$ represents known input variables, and $U=T\cup O$ represents unknown variables including the non-constant variables including temporal variables $T$ and output variables $O$. $E$ includes all the constraints of the circuit $C(\mathcal{X})$.
3 Problem Statements
Detecting overconstrained or underconstrained circuits in Circom boils down to an algebraic inquiry into the number of solutions for a quadratic equations system over a finite field. This inherent connection naturally leads us to model Circom circuits as such systems and leverage algebraic methods for analysis. Moreover, the design principle of R1CS forces Circom constraints under quadratic computation, resulting a maximum degree of 2 for the equivalent equations. Our method follows the classification-and-discussion scheme. Our strategy categorizes circuits by degree and applies specialized solutions within each category, leading to significant efficiency and scalability gains.
3.1 Constraint Problems
The construction of arithmetic circuits is plagued by two critical challenges: overconstrained and underconstrained circuits. Programmer-authored circuits, even from seasoned ZK developers, often bear inadvertently miscalibrated constraints, leading to these issues. While completely eradicating these issues remains a persistent hurdle, current approaches fall short of effectively addressing them. This subsection tackles this double-edged sword head-on, proposing a novel approach to ensure the construction of robust and reliable arithmetic circuits.
Overconstrained bugs. As described in the work by Labs et al. [27], overconstrained circuits are unable to uphold the completeness of zero-knowledge proof. A Circom program is deemed overconstrained when its corresponding polynomial constraints are overconstrained, indicating that the circuit cannot produce any valid outputs regardless of the inputs provided. This inconsistency in the constraint system for a given input characterizes the state of being overconstrained. Despite their deceptive simplicity in concept and the misconception that they are less critical than other types of attacks, denial of service (DoS) attacks are often underappreciated regarding their security implications. Notably, DoS attacks require neither high skill nor deep knowledge,, contributing to their low cost for malicious attackers as highlighted by Zcash [30].
Underconstrained bugs. Recent studies have discovered that underconstrained bugs in arithmetic circuits can lead to various security threats, including the ability for attackers to forge signatures, steal user concurrency, or create counterfeit crypto-coins. A Circom program is classified as underconstrained when the corresponding polynomial constraints are also underconstrained. This condition arises when there exists an input $in$ for which multiple distinct outputs $out_{1},out_{2},\ldots$ can be produced, and the circuits $C(in,out_{1}),C(in,out_{2}),\ldots$ are all considered valid. In essence, an underconstrained circuit program may unintentionally permit the acceptance of unwanted signals, making it susceptible to exploitation by malicious entities.
3.2 Categories of Elements
In our method, categorizations are necessary to account for the various types of variables present in arithmetic circuits. Some variables are determined by external inputs, while others depend on other variables, necessitating the need to classify them accordingly. Aligning with Circom’s degree-2 limit, it is pertinent to categorize the variables based on the degree. Furthermore, due to the computational complexity of verifying certain circuits, we are limited to providing coarse-grained results, thereby highlighting the importance of categorizing the outcomes as well.
Category of Variables. The variables in a Circom circuit consist of the input signals $I$, temporary signals $T$, and output signals $O$. The user provides the values for $I$, which can be regarded as known values $K$. On the other hand, the variables in $O\cup T$, whose values are not constant, are regarded as unknown variables $U$ within the constraints. The key issue of the constraint problem is identifying the unique characteristics of the $O$ variables.
Category of Equations and Circuits. The constraint $e$ is an algebraic polynomial equation whose degree is under 2 over a finite field $F$, while the circuits are a system of algebraic polynomial equations whose degree is under 2 over a finite field, with their generator variables being $U$.
The three types of equations (constraints) to consider are precise linear equations, $K$-coefficient linear equations, and quadratic equations. A precise linear equation is defined by having linear generator variables with constant coefficients. Conversely, a $K$-coefficient linear equation includes at least one generator variable with a coefficient represented by the known variables $K$, and none of the generator variables’ coefficients are represented by the unknown variables $U$. Lastly, a quadratic equation is distinguished by containing at least one generator variable with a coefficient represented by the unknown variables $U$.
The system of equations, namely circuits, can be categorized into three types based on the constraints outlined above: precise linear circuits, $K$-coefficient circuits, and quadratic circuits. A precise linear circuit consists solely of precise linear equations, while a $K$-coefficient circuit must contain at least one $K$-coefficient linear equation and no quadratic equations. Finally, a quadratic circuit should include at least one quadratic equation.
Category of Results. As shown in Fig. 1, $\text{AC}^{4}$ yields six result types: algebraic overconstrained, precisely overconstrained, algebraic exactconstrained, precisely exactconstrained, precisely underconstrained, and unknown. The algebraic overconstrained and algebraic exactconstrained results are collectively referred to as algebraically solved results, while precisely overconstrained and precisely exactconstrained results are collectively referred to as precisely solved results.
The verification problem is refined to create a more detailed classification system for the output, based on constraints safety metrics. As for the unknown cases that are not fully solved by the solver, they are categorized into two intermediate cases separated by constraint stress. The two intermediate cases, algebraically exact constrained and algebraically over constrained, provided more detailed information for analysis. Besides this category, this method unlocks partial information about unsolvable circuits and speeds things up.
Fig. 1 also provides a cross-reference between our result category and the typical ground truth of constraints. Among those 5 practical output types except unknown  3 precise categories of them: precisely underconstrained, precisely exactconstrained, and precisely overconstrained, strictly maps into the underconstrained, exactconstrained, overconstrained respectively. The 2 medium status algebraic exactconstrained and algebraic overconstrained, caused by potential reduced-rank phenomenon while simplifying the equation matrix in some cases. Our approach cannot ascertain equation system solution counts, reflected in matrix rank. Cases labeled with algebraic exactconstrained may lie in underconstrained or exactconstrained of ground truth, while cases labeled with algebraic overconstrained occur in exactconstrained or overconstrained.
Algebraically overconstrained indicates no algebraic solutions, but special inputs can shift results to exactconstrained or underconstrained. The circuit is algebraically exactconstrained, indicating it has exactly one solution algebraically. However, Special inputs can trigger underconstrained states.
Unknown results may stem from circuit size, computational complexity, or the inherent NP-hardness of solving quadratic equations over finite fields.
By combining the algebraic exactconstrained and precisely exactconstrained into a synthesized state, namely general exactconstrained, cases in this state are never overconstrained, though they could fall in underconstrained. Similarly, the general overconstrained state provides a guarantee for cases to avoid underconstrained state. We introduce a relaxed constraint property metric that efficiently filters out under/overconstrained cases.
Figure 1: The category of results from AC4.
3.3 Problem Reduction
This section demonstrates the reduction of the under/overconstrained problems of a circuit to the number of solutions of the corresponding polynomial equations.
Definition 4 (Polynomials over finite fields)
For a finite field $\mathbb{F}$ and formal variables $X_{1},\ldots,X_{k}$, $\mathbb{F}[X_{1},\ldots,X_{k}]$ denotes the set of polynomials in $X_{1},\ldots,X_{k}$ with coefficients in $\mathbb{F}$. By taking the variables to be in $\mathbb{F}$, a polynomial $f\in\mathbb{F}[X_{1},\ldots,X_{k}]$ can be viewed as a function from $\mathbb{F}^{k}\rightarrow\mathbb{F}$.
To reduce the circuit constraint problems to the number of solutions of polynomials, we have the following proposition:
Proposition 1 (Circuit-Poly Reduction)
For each circuit $A=\langle C(X),e\rangle\in\langle C(\mathcal{X}),E\rangle$, there exists $A^{\prime}\in\mathbb{F}[X_{1},\ldots,X_{k}]$, such that $A^{\prime}=\mathbb{M}(A)$ , where $\mathbb{M}$ is a bijection function from a circuit to quadratic polynomials over finite fields.
[…]
Based on the argument above, completely checking the constraint status of circuits can be achieved by solving polynomials. More concretely, the constraint problem corresponds to figuring out the number of solutions for polynomials.
4 Methodology
4.1 AC4 Framework
Fig. 2 illustrates the overall workflow of our proposed tool, $\text{AC}^{4}$, which receives a Circom program file as the input, and compiles the program in debug mode into an R1CS format file along with a sym file. To retrieve the variables and constraints from the compiled files, a task-specific parser is implemented to handle those compiled files, focusing on unknown variables within constraints. After that, the extracted constraints are then simplified into equations. A categorizing mechanism is leveraged in Section 3.2 to determine the specific category of the circuit according to its simplified equation system. Based on the category result of circuits, three types of solvers is employed to address them respectively.
Figure 2: The framework of AC4.
In the context of a precisely pure linear circuit, the system of linear equations can be represented in matrix form. While it may seem natural to rely on rank to determine the number of solutions, it is more important to prioritize verifying the uniqueness of the output variable set. Therefore, solving the system demands a suitable method for linear equations. Specifically, in $\text{AC}^{4}$, the Gauss-Jordan method[42] is utilized to obtain the solution of the circuits as well as to ascertain the uniqueness of the set of output variables.
The system of $K$-coefficient linear equations corresponding to an algebraically linear circuit can be represented in matrix form. However, it is imperative to note that algebraic computation does not inherently account for the division by zero problem. Consequently, for certain specific inputs $I$, the rank of the matrix may decrease, potentially altering the outcome of $\text{AC}^{4}$. Rank reduction risks are mitigated by strategically identifying and treating special input cases. Our approach addresses this challenge through heuristic approaches to improve the efficiency and accuracy.
For quadratic circuits, we solve the coefficients and verify if the circuit will be underconstrained with specific inputs. Then, we use $\text{AC}^{4}$ to directly solve the systems of quadratic equations with Gröebner basis[25] and check the uniqueness of the output variables.
Here, an example will be used to demonstrate our workflow and enhance readers’ comprehension. The template in Fig. 3 refers to the Circom circuit Decoder, which converts a w-ary number into a one-hot encoded format.
1template Decoder(w) {
signal input inp;
3 signal output out[w], success;
var lc=0;
5 for (var i=0; i < w; i++) {
out[i] <– (inp == i) ? 1 : 0;
7 out[i] * (inp-i) $=$ 0;
lc = lc + out[i];}
9 lc ==> success;
success * (success -1) $=$ 0;
11}
component main = Decoder(2);
Figure 3: Decoder(2)
Upon setting $w=2$, the Decoder(2) emerges as a 2-bit decoder. This circuit is then transformed into a format denoted as $C(K,U,O,E$, where the sets are defined as follows: $K=\{inp\}$, $U=\{out_{1},out_{2},success\}$, $O=\{out_{1},out_{2},success\}$, and $E=\{out_{0}\times inp=0,out_{1}\times(inp-1)=0,success\times(success-1)=0\}$.
Based on the classification of circuits mentioned above, it is evident that this circuit is quadratic. We begin by treating the variable $inp$ as a coefficient and proceed to solve the undetermined coefficients of each equation. This process yields the solutions [$[\{inp=0\},\{inp=1\}]$. Upon substituting $\{inp=0\}$ into $E$, it results in $E=\{0=0,out_{1}\times(inp-1)=0,success\times(success-1)=0\}$. Notably, it is evident that the $out_{0}$ can take on any value over the finite field, signifying that the circuit is underconstrained.
4.2 Checking Circuits: Category-Based Algorithms
From above, we divide the overall circuit verification problem space into three types: precisely linear, $k$-coefficient linear, and quadratic separately. In this section, we introduce those algorithms in detail.
Precisely linear. The majority of circuits are precisely linear, so they can be represented by a coefficient matrix $A$ and a constant vector $b$. While it is commonly thought to use the rank of the equation matrix to determine the number of solutions, the solver’s objective is only to verify the uniqueness of output variables. In contrast, rank-reducing methods impose restrictions on the uniqueness of all variables, including temporary variables. Using the Gauss-Jordan method to get the rank of the corresponding matrix is analogous to getting the solution of the system of equations. Consequently, there is no significant difference in the time complexity. The pesudo-code is given in Algorithm 1.
Algorithm 1 Algorithm for checking linear circuits
1:procedure CheckLinearCircuit($\langle C(\mathcal{X}),E\rangle$)
2:     $A,b\leftarrow\textsc{LinearEquationsToMatrix}(E,U)$triangleright $A$ is the coefficient matrix and $b$ is the constant matrix
[…]
4:     if IsEmpty(sols) then
5:         return precisely overconstrained
6:     else if CheckUniqueness(sols, $O$ then
7:         return precisely exactconstrained
8:     else
9:         return underconstrained
10:     end if
11:end procedure
Qudratic. A heuristic approach will be used to identify input variables that may cause the circuit to become underconstrained (shown in Algorithm2), aiming to improve accuracy and efficiency. The coefficients of the unknown variables are treated as equations and their solutions will be stored in a frequency table for counting the occurrences of each solution. Then, the solver sorts the solutions based on their frequencies and iteratively substitutes them into the original system of equations to obtain a new system, which is then solved to check if the circuit is underconstrained. If the input variable causing the circuit to become underconstrained cannot be identified, the system of equations will be directly solved with Gröebner basis to verify the constrained situation of the circuit. The pseudo-code of checking quadratic circuits is given in Algorithm3.
Algorithm 2 check circuits by the coefficients of the output variables
procedure GetUndeterminedCoeffSols($K,U,E$)
     $UCSolCountTable\leftarrow EmptyHashTable$
     for all $eq\in E$ do
         $sols\leftarrow\textsc{SolveUndeterminedCoefficientEq}(E,K,U)$
         if $sols\neq\emptyset$ then
              $UCSolCountTable[sol]++$
         end if
     end for
     return $UCSolCountTable$
end procedure
Algorithm 3 Algorithm for checking quadratic circuits
procedure CheckQuadraticCircuit($\langle C(\mathcal{X}),E\rangle$)
     do the same thing to undetermined coefficients and check underconstrained
[…]
     if IsEmpty(sols) then
         return algebraically overconstrained
     else if CheckUniqueness(sols, $O$ then
         return algebraically exactconstrained
     else if $\textsc{Size}(sols)>1$ then
         return underconstrained
     else
         return unknown
     end if
end procedure
$K$-coefficient linear. Although quadratic in nature, these equations are considered linear due to their quadratic terms containing one or more known variables, Hence, they are treated as linear circuits and represented in matrix form to enhance verification efficiency. For these circuits, we still use the previous heuristic approach given in Algorithm 2, to determine if there are special inputs that can cause the underconstrained condition, aiming to enhance efficiency. Afterwards, a $k$-coefficient linear circuit can be succinctly represented by a coefficient matrix $A$ and a constant vector $b$, and checked as a linear case. The pseudo-code of checking $k$-coefficient circuits is given in Algorithm4.
Algorithm 4 Algorithm for checking $k$-coefficient circuits
1:procedure CheckKCoeffcientCircuit($\langle C(\mathcal{X}),E\rangle$)
2:     $udcSols\leftarrow\textsc{GetUndeterminedCoeffSols}(K,U,E)$
3:     for all $sol\in udcSols$ do
[…]
5:         if Check $C(U,K,O,E_{sub})$ is underconstrained then
6:              return underconstrained
7:         else
8:              continue
9:         end if
10:     end for
11:     $A,b\leftarrow\textsc{LinearEquationsToMatrix}(E,U)$
[…]
13:     if IsEmpty(sols) then
14:         return algebraically overconstrained
15:     else if CheckUniqueness(sols, $O$) then
16:         return algebraically exactconstrained
17:     else if $\textsc{Size}(sols)>1$ then
18:         return underconstrained
19:     else
20:         return unknown
21:     end if
22:end procedure
4.3 Complexity Analysis
This section discusses the time complexity of each category. First, for linear or part of $K$-coefficient linear circuits, the Gauss-Jordan solution yields a complexity of $O(n^{3})$ [2]. As we discovered that the majority of cases are precisely linear circuits based on a statistical analysis of the Circomlib [23], We expect our algebraic-based method to introduce a large improvement in performance compared with SMT solvers, as SMT solvers tend to convert constraints into Conjunctive Normal Form (CNF) which ignores potential mathematical transformation and simplification within equation systems. When considering a quadratic equations system over a finite domain, the time complexity is proven to be NP-hard [17]. Even with complex cases, quadratic solvers offer a practical approach.
4.4 Optimization and Implementation
Pre-implementation. Before solving the equations, it is crucial to first examine whether there are any output variables that are not present in any constraint, as the presence of such a variable indicates that the circuit is underconstrained. Once identified, these variables can then be addressed appropriately. Furthermore, it is important to note the existence of equations with no unknown terms. These equations should be filtered out as they do not contribute to the determination of the output variables’ uniqueness. In addition, it is necessary to carefully assess the type of each equation to establish the type of the circuit being analyzed ultimately. Moreover, in scenarios involving $k$-coefficients, it is possible for the result to be algebraically overconstrained, thereby requiring a methodical approach to narrow down its range. To achieve this, the first step involves checking for linear input variables and treating them as unknown in order to attain more accurate solutions. This systematic process enhances the precision and reliability of the solutions obtained.
Binary Optimization. We improved the efficiency of checking the binary circuits by first encoding them into binary format over a finite field. Then, we implemented a detector to verify whether a quadratic circuit is a binary circuit by ensuring that all unknown variables adhere to the form $u\times(u-1)=0$. This binary checking process was employed to validate the circuits further.
Handling of finite field operations. In our experiments, we observed that only a small number of binary circuits would experience overflow in the finite domain. Despite the availability of relatively efficient algorithms for polynomial operations over finite fields, such as Berlekamp’s algorithm[10], the large characteristic of the finite field of the circom can result in less efficient operations. As a result, we opted to employ integers in the practical implementation. Additionally, we utilized a table to strategically map overflow values and their corresponding original values within a select few binary circuit programs. This approach was implemented to enhance the overall efficiency of the system.
5 Evaluations
Environment configuration. The evaluations were performed on a machine with a 16-core AMD Ryzen 9 5950x processor and 128GB of DRAM.
Check Benchmark. We settle a representative benchmark to evaluate the effectiveness and accuracy of $\text{AC}^{4}$ and Picus [37]. The benchmark was based on the data set provided by $\text{AC}^{4}$ and Picus, which had been collected from Circomlib, the standard library from Circom. All cases of our benchmark were divided into two parts: Circomlib-utils benchmarks, which was extracted from the Circomlib utils library and had a smaller size, and Circomlib-core benchmarks, which was extracted from the Circomlib core library and had a larger size [37].
5.1 The Conditions of the Data Set
The size condition Table 1 presents the size condition of the benchmark set. Notably, within the benchmark set Core, there are three circuits with over 200,000 constraints, which are believed to be unsolvable using standard computers. Interestingly, in these three cases, only Sha256compression results in an out-of-memory exception due to the presence of around 20,000 unknown variables, resulting in the creation of a very large matrix. In contrast, the other two cases each have only one known variable, making the circuits easier to check.
Table 1: The size condition of the benchmark sets
Benchmark set # of circuits Avg. # of expr_num Avg. # of unknown_num Avg. # of out_num
utils 66 1992 12 9
core 110 6941 200 30
total 176 5085 130 22
We categorize circuits based on their constraint numbers in Table 2, considering those with a constraint number less than 100 as ”small” circuits, those with a constraint number in the range $[100,1000)$ as ”medium” circuits, and those with a constraint number in the range $[1000,\infty)$ as ”large” circuits. This classification, which may be rudimentary, does not consider output signals. It is evident that small circuits can be solved directly, whereas other circuits may require specific restrictions to be solved within an acceptable time frame.
Table 2: The size category condition of the circuits
Benchmark set utils core total
small 49 63 112
medium 8 24 32
large 9 23 32
The type condition As shown in the table 4, The benchmark set comprises 176 circuits, distributed as 71.6% precisely linear, 6.2% $K$-coefficient linear, and 22.2% quadratic, mirroring the range of circuit types encountered in real-world applications. Upon further examination of the $K$-coefficient and quadratic cases, it becomes evident that a significant portion of them are binary circuits associated with cryptographic algorithms. Instead of exhaustive arithmetic solutions, focusing on binary optimization strategies offers promising efficiency gains in checking these circuits. For instance, the circuit Num2BitsNeg@bitify_256 is a quadratic circuit designed to convert an integer over a finite field into a 256 bit-vector, with all its unknown variables constrained by $o\times(o-1)=0$, indicating that they are strictly bits with their value limited to 0 or 1. Appropriately identifying these circuits and applying binary methods for solving them represents a more efficient approach compared to arithmetic methods.
5.2 Benchmark Metrics
Checked Ratio. The solved benchmarks are initially categorized into three types: precisely solved, algebraically solved, and unknown. The precisely solved cases are further categorized into three types: precisely exactconstrained, which means that the circuit is safe; precisely overconstrained, indicating that the circuit is overconstrained; and precisely underconstrained, which suggests that the circuit is underconstrained. We omit this metric as $PS$ (Precisely Solved).
On the other hand, we propose a more relaxed metric, $AS$ (Algebraic Solved), by remapping cases that failed to obtain precise results into two new categories: algebraic exactconstrained and algebraic overconstrained as introduced in Section 3.2, to reject overconstrained and underconstrained respectively. Note that $\text{AC}^{4}$ still returns unknown for extremely computationally exhausting cases.
Checking time. In this metric, our emphasis lies on the benchmark data on the parts that both $\text{AC}^{4}$ and Picus can solve, and then proceed to compare the checking time of the two tools.
5.3 Benchmark Results
Checking Time The checking time of the benchmark set is presented in Table 3, with a focus on the relationship between benchmark size and solving difficulty. It is observed that the larger benchmarks present greater difficulty in solving, resulting in longer solution times. However, in some large circuits, despite a high number of constraints, the number of output signals is limited to 1, allowing for a relatively easy check once the file read operation concludes. Nonetheless, there exist extremely large circuits that are characterized by a linear type, yet their extensive constraint size renders the Gaussian-Jordan method challenging to execute, leading to out-of-memory issues. Conversely, smaller cases are relatively straightforward to solve directly but due to limitations, the precise solution is not attainable, resulting in algebraic outcomes.
Fortunately, the majority of practical Circom circuits are of the precisely linear type. This characteristic is advantageous because medium or large circuits that are not precisely linear cannot be solved directly due to unknown outcomes. Therefore, alternative methods must be utilized to bypass the direct solution of non-linear circuits. Additionally, providing a more accurate range to determine whether to solve the circuits can help avoid obtaining algebraically solved results. This approach is crucial for maintaining clarity and precision in circuit analysis.
Table 3: The efficiency benchmark of $AC^{4}$ is categorized by size and includes the time for precisely solved and algebraically solved results. The terms $K,U,O,v,PS,AS$ represent known, unknown, output, variables, precisely solved, and algebraically solved, respectively. The explanation of $PS,AS$ is discussed in Section 5.2.
Benchmark Utils Core overall
Size small medium large overall small medium large overall
Avg. # variable K 15 424 14309 2014 36 435 24984 5159 3973
U 4 18 58 13 11 78 26 29 23
O 2 2 57 9 10 78 26 29 21
total v 18 442 14366 2026 47 512 25010 5188 3995
constraints 14 244 14230 1993 17 474 24923 5144 3956
Avg. PS Time(s) 0.57 0.52 9.06 1.83 0.42 2.4 14.59 3.71 3.03
Avg. AS&PS Time(s) 1.26 2.45 8.33 2.41 0.40 2.4 13.47 3.48 3.06
Checked Ratio. The checked ratio of the benchmark sets categorized by circuit type is presented in Table 4. The data indicates a significantly high solved ratio for precisely linear circuits. However, Substantial constraints in circuits result in very large augmented matrices. As a consequence, the solve ratio of $k-$coefficient circuits is relatively low. This is attributed to the adoption of an algebraic strategy for $k-$coefficient circuits, which in turn leads to the direct solving of small-size $k-$coefficient circuits without returning a precise result.
Heuristic algorithms initially assess potential underconstraint in quadratic circuits. We then apply classic algorithms for solving quadratic equations over finite fields to obtain direct solutions. Moreover, in the case of specific binary circuits such as BinSum and Num2Bits, a detection method is utilized to verify the boolean nature of unknown variables. Upon confirmation, a specialized method for solving binary circuits is employed in lieu of arithmetic-based approaches. Most circuits are thoroughly verified before processing.
Table 4: Main Experiment results, we categorize circuits based on their types: pl means precisely linear type, kcl means k coefficient linear, q means quadratic, ps means precisely solved, as means algebraically solved.
Benchmark Utils Core overall
type pl kcl q overall pl kcl q overall
ps_num 45 3 7 55 75 3 23 101 156
as_num 3 3 5 11 2 2 4 8 19
total_num 48 6 12 66 77 5 27 109 175
ps_rat 94% 60% 58% 83% 96% 60% 85% 91% 89%
ps&as_rat 100% 100% 100% 100% 98% 100% 100% 99% 99%
5.4 Comparison with Baseline
In the paper [37], Picus, an implementation of the tool $\textbf{QED}^{2}$, is utilized for verifying the uniqueness property (under-constrained signals) of ZKP circuits. This implementation employs static program analysis and SMT solvers checking to achieve the verification of the uniqueness of ZKP circuits. Subsequently, we benchmarked the Picus program using the same benchmark set.
The completed Picus ratio for the utils benchmark set is 80.33%, for the core benchmark set it is 62.61%, and for both sets it is 69.36%. Additionally, the average solving time for the Picus solver is 425.73 seconds for the utils benchmark set, 317.88 seconds for the core benchmark set, and 365.52 seconds for both sets.
(a) Comparison between the checking time of $\text{AC}^{4}$ and Picus.
(b) Comparison between checked ratio of $\text{AC}^{4}$ and Picus
Figure 4: The checking time and checked ratio comparisons between $\text{AC}^{4}$ and Picus
The comparison of the two tools based on the solved ratio can be found in Fig. 4b. Our advantage lies in precisely linear cases as well as some heuristic methods for $k-$coefficient and binary cases.
Fig. 4b shows the solving time comparison of the two tools. Our advantage lies in solving problems arithmetically rather than encoding them into SMT formulas, making our tools much faster.
5.5 Analysis
Table 3 illustrates a significant correlation between the size of the circuit and the time required for checking. Despite their size, large circuits keep checking times low due to two critical factors. Firstly, the predominance of linear circuits, efficiently addressed by the Gauss-Jordan solver, contributes to the efficiency in solving a majority of cases. Additionally, optimizations targeting binary cases within k-coefficient and quadratic circuits serve to further reduce the solving time despite their slightly larger size.
The checked ratio is high for precisely linear cases; however, there are some cases for which we cannot provide precisely checked results due to the lack of an algebraic solution. Special input scenarios can reduce the augmented matrix’s rank, yielding a solution. In contrast, for the $K$-coefficient cases, the checked ratio is relatively low, even when a mixed solver is employed. This is because we did not implement certain static analysis methods to elaborate the benchmark set, which would have facilitated the checking of circuits. Conversely, the checked ratio is relatively high for quadratic cases, as most of these cases are in binary form. The detection method and optimization for binary quadratic circuits have significantly enhanced efficiency in checking these cases.
AS Analysis. We point out that the cases that occur in the AS metric are worth discussing. In Core data split, cases that cannot be precisely solved consist of nearly 7% of the total amount, while this number increases to 17% in Utils data split. Our proposed AS mechanism enables our tool to investigate more detailed situations of constraints. By excluding an incorrect option among underconstrained and overconstrained, we narrow the possible status of these cases. Experimental results demonstrate that our tool offers valuable hints to almost 100% cases in both Core and Utils splits.
Failure Analysis. Picus utilized their algorithm Uniqueness Constraint Propagation to solve all the $k-$coefficient circuits. The algorithm simplifies verification/resolution by strategically assigning values to unknown variables through partial equation analysis, reducing unknowns. However, our methods forgo static analysis, limiting $K$-coefficient case checking.
6 Related Work
Verification on ZK programs. Circomspect [12] serves as a static analyzer and linter designed for the Circom programming language. Picus [37] employs a static analysis method combined with an SMT solver. Circomspect and Picus share certain similarities with our work. However, Circomspect can only annotate a portion of underconstrained variables. Picus is immune to this issue but remains a low efficiency due to the extensive use of the SMT solver.
ZK circuit formalization has seen notable advancements: Shi et al. [41] proposed a data-flow-based algorithm for R1CS standardization, to enhance performance across R1CS tasks. Wen et al. [47] formalized and classified Circom vulnerabilities, introducing a CDG-based static analysis framework. CODA [29] emerged as a statically typed language for formal ZK system property specification and verification. Soureshjani et al.[43] applied abstract interpretation to examine Halo2 circuit properties, on the Plonkish matrix using an SMT solver.
Research on ZK compilers with theorem proving is ongoing. Bangerter and Almeida et al.[5, 4, 1, 6], pioneered verifiable compilation using Isabelle/HOL [36]. PinocchioQ [20] verified Pinocchio[38], a C-based zk-SNARK, with Compcert [28] based on Coq [11]. Chin and Coglio et al [15, 16]. employed ACL2 [26] on Leo, while Bailey et al. [3] established formal proofs for six zk-SNARK constructions using the Lean theorem prover [34].
Algebra computation in circuit verification. Algebraic computation is a long-standing but niche endeavor in verification. Watanabe [46] utilize Gröbner Basis [19, 18] and polynomial reduction techniques to verify arithmetic circuits in computing and signal processing systems. Lv et al. [32, 31] used a computer-algebra-based approach for the formal verification of hardware. Scholl et al. [40] combined Computer Algebra with Boolean SAT solver to verify divider circuits.
However, we emphasize that most of the work in this field actively reduces program or hardware problems to algebra. To the best of our knowledge, we are the inaugural contributors to the application of algebraic methods in ZKPs.
7 Conclusion
Using an algebraic computation system, we have developed a technique for detecting zero-knowledge bugs in arithmetic circuits. This approach is designed to address issues caused by underconstrained or overconstrained arithmetic circuits. Unlike the traditional method of encoding the problem into formulas and utilizing an SMT solver, our approach leverages arithmetic computations, significantly improving efficiency. Additionally, to enhance the verification process, we have incorporated heuristic methods. Notably, our method directly analyzes arithmetic circuits and does not depend on a specific domain-specific language (DSL), making it applicable to a wide range of DSLs that support zk-SNARKs. Our implementation utilizes the open-source algebra computation system, SymPy, in a tool called $\text{AC}^{4}$. We conducted an evaluation involving 176 Circom circuits, and our approach successfully verified approximately 89% of the benchmark sets within a short timeframe.
In order to enhance the elaboration of the data set prior to direct solution, our approach will encompass adopting methods such as static program analysis. This will enable us to acquire comprehensive information about the circuits before applying our methods. Additionally, we aim to conduct extensive evaluations of commonly utilized circuits and design specialized algorithms for them. Furthermore, the use of theorem proving will be employed to demonstrate the correctness of a series of specific circuits. Ultimately, our goal is to develop or modify existing open-source computer algebra tools to tailor a solver for determining the number of solutions to a system of polynomial equations over a finite field.